ELECTRONICALLY EXCITED STATE DYNAMICS OF THE ALKOXY RADICALS

\underline{SANDHYA GOPALAKRISHNAN}, LILY ZU, TERRY A. MILLER, The Ohio State University, Dept. of Chemistry, Laser Spectroscopy Facility, 120 W. 18th Avenue, Columbus, Ohio 43210.

We have measured the lifetimes of some of the prominent bands observed in the moderate-resolution, jet-cooled laser induced fluorescence excitation spectra of the B̃ ← X̃ transition of several alkoxy radicals. High-resolution rotationally resolved studies have given isomer and conformer specific assignments for all the bands for which lifetimes have been measured. We report lifetime trends observed as a function of isomer and conformer, as well as for level of vibrational excitation. The implications of these observations for the dynamics of the B̃ state will be discussed.